OpenStudy (anonymous):

rewrite sec∅(1-sin^2∅) in terms of cos∅ and simplify

OpenStudy (anonymous):

\[\sec ∅ (1-\sin^2∅)\]

OpenStudy (anonymous):

\[\sec(\theta) = 1/\cos(\theta)\]\[1-\sin^2(\theta) = \cos^2(\theta)\]\[\cos^2(\theta)*(1/\cos(\theta)) = \cos(\theta)\]